OpenStudy (anonymous):

"Find a polynomial P(x) having real coefficients, with the degree and zeroes indicated. (Assume the lead coefficient is 1.)" No idea what it means by the "assume" part. I think what I have to do is basically turn these numbers into an equation...? degree 4; x = -1, x = 1 + 2i

myininaya (myininaya):

If x=-1 is a zero then x-(-1) is a factor of the polynomial If x=1+2i is a zero then x-(1+2i) is a factor of the polynomial Ok now we want real coefficients so what does that tell us another zero would be?

OpenStudy (anonymous):

So, (x + 1) is one, but don't we have to change that 2i? On my other problems, it came out to (x^2 + 4) (of course, I'm sure I might doing it wrong...) But then, what happens to the 1?

myininaya (myininaya):

Ok so if i tell you -2+2i is a zero, then -2-2i is also a zero. Since (x-(-2+2i))(x-(-2-2i))=(x+2-2i)(x+2+2i)=([x+2]-2i)([x+2]+2i) =(x+2)^2+2i(x+2)-2i(x+2)-4i^2=(x+2)^2-4(-1)=(x+2)^2+4 <see no imaginary parts anymore that is what we want for this problem So seeing this example fill in the below blank: For this problem you say if 1+2i is a zero, then __???____ is also a zero

myininaya (myininaya):

Another example: If 2+i is a zero, then 2-i is also a zero SINCE (x-(2+i)(x-(2-i))=(x-2-i)(x-2+i)=([x-2]-i)([x-2]+i)=(x-2)^2+i(x-2)-i(x-2)-i^2=(x-2)^2-(-1)=(x-2)^2+1 see no imaginary parts anymore

myininaya (myininaya):

Also I'm asking you to do is give the conjugate of 1+2i Do you know what it is?

myininaya (myininaya):

All* not also sorry

OpenStudy (anonymous):

I'm sorry, but I can't follow you at all. I'm still trying to work out the string of equations you gave me. Also, what do you mean by "conjugate"? I'm sorry I'm being so difficult.

myininaya (myininaya):

The conjugate of a+bi is a-bi The conjugate of a-bi is a+bi The conjugate of 2+i is 2-i The conjugate of -2+2i is -2-2i The conjugate of -2-2i is -2+2i The conjugate of 2-i is 2+i The conjugate of 41-3i is 41+3i and so on...

OpenStudy (anonymous):

Ah, so, if x = 1 + 2i, it also equals 1 - 2i. Is that it? So...there's x = 1+ 2i, 1 - 2i, and -1. But since it's to the fourth degree, does this mean -1 repeats?

myininaya (myininaya):

no 1+2i is not equal to 1-2i I'm saying if one zero is x=1+2i then another zero is x=1-2i Since x=1+2i => x-(1+2i) is a factor x=1-2i => x-(1-2i) is a factor And when you multiply these you get real coefficients like this (x-(1+2i))(x-(1-2i)) (x-1-2i)(x-1+2i) ([x-1]-2i)([x-1]+2i) [x-1]^2+2i[x-1]-2i[x-1]-4i^2 [x-1]^2+0-4(-1) [x-1]^2-4(-1) [x-1]^2+4 See we have real coefficients

OpenStudy (anonymous):

"Equals" was bad word choice on my part, I meant to ask if it was another zero, like -1 and 1 + 2i. My bad. I know it's taking me awhile. I'm trying to see if I can work this out better...

myininaya (myininaya):

\[(x-(-1))(x-(1-2i))(x-(1+2i))(x-a)\] I said x=a is another zero since it was a 4th degree polynomial x=a is a zero => x-a is a factor of the polynomial now we can clean this up a bit we have \[(x+1)((x-1)^2+4)(x-a)=(x+1)(x^2-2x+1+4)(x-a) \] \[=(x+1)(x^2-2x+5)(x-a)\] I will let you choose a and multiply the rest out

myininaya (myininaya):

I think a=1 or a=0 is a good choice

myininaya (myininaya):

but you can choose whatever

OpenStudy (phi):

Though the question does not strictly rule out a fourth root, I would assume they want a repeated root x=-1 to exactly meet their criteria

myininaya (myininaya):

Well then that would mean they should have also said x=1-2i is a zero

OpenStudy (phi):

They did when they said all the coeffs are real

myininaya (myininaya):

but they didn't say x=-1 multiplicity 2

myininaya (myininaya):

so why does it have to be x=-1 mult. 2

OpenStudy (phi):

no, but they did want a 4th degree, with the given (and implied) zeros

OpenStudy (phi):

and it's not wrong to use a repeated root...

myininaya (myininaya):

yeah but like i said they didn't list x=1-2i as a zero but the question implies that that should be a zero

myininaya (myininaya):

but the question didn't say x=-1 mult. 2

myininaya (myininaya):

it said find a polynomial it says "a" meaning that could be more than one answer

myininaya (myininaya):

depending on what you let a equal

OpenStudy (anonymous):

Okay, it took me awhile, but I think I got them all worked out... \[(x+1)(x^2-2x+5)(x-1) = x^4 - 2x^3 + 4x^2 + 2x - 5 \] \[(x+1)(x^2-2x+5)(x-0) = x^4 - x^3 + 3x^2 + 5x\] \[(x+1)(x^2-2x+5)(x+1) = x^4 + 2x^2 + 8x + 5\]

myininaya (myininaya):

lol so which one are you gonna go with?

myininaya (myininaya):

phi was suggesting you go with the bottom one it makes no difference to me based on the wording of the question

OpenStudy (anonymous):

...I honestly have no idea. It looks like they could all be correct, but I have no way of knowing which one my professor will take. My book actually does mention something about repeating roots, but it's pretty vague about it. So, to play it safe: Turn in the one-question page sitting in front of me and hope for the best. Heh. Thank you so much for helping me! I know it took forever, but I understand this a lot better now!

myininaya (myininaya):

That is good :)